OpenStudy (anonymous):

Find the solution to the differential equation: (dy/dx) + (y/3)= 0, y(0)=10

OpenStudy (anonymous):

seperate the equation with the x's on one side and the y's on the other and then take the derivative of both sides

OpenStudy (anonymous):

Can someone help me start it? I'll be able to get it once I get it seperated, but I'm having a hard time starting.

OpenStudy (anonymous):

I lost the equation. What was it?

OpenStudy (anonymous):

(dy/dx) + (y/3)= 0, y(0)=10

OpenStudy (anonymous):

ok. sorry. It is an integral; not a derivative. dy/dx=-y/3 3 dy= -y dx 3/-ydy = 1 dx then integrate both sides...

OpenStudy (anonymous):

then the initial condition y(0)=10 determines the constant

OpenStudy (anonymous):

Thank you, I think I got confused by the y(0)= 10.